OpenStudy (anonymous):

A segment with endpoints A (4, 2) and C (1, 5) is partitioned by a point B such that AB and BC form a 1:3 ratio. Find B. a. (1, 2.5) b. (2.5, 3.5) c. (3.25, 2.75) d. (3.75, 4.5)

Directrix (directrix):

AB/BC = 1/3 So, BC = 3*AB